A 72N force accelerates mass (m1) by 5.0 m/s2. The same force acts on another mass (m2) and imparts to it an acceleration of 20.0 m/s2. What accelerates will the same force impart to the two masses when they are joined together?
1
Expert's answer
2020-08-31T12:50:13-0400

According to newton's second law;

"F=ma"

Therefore; "M1=\frac{F}{a}=\frac{72}{5}=14.4kg"


"M_2=\frac{5}{20}M_1=\frac{5}{20}\times14.4=3.6kg"


"a=\frac{F}{(M_1+M_2)}=\frac{72}{(14.4+3.6)}=4m/s^2"